PHOTODISSOCIATION DYNAMICS OF THE PHENYL RADICAL VIA PHOTOFRAGMENT TRANSLA-TIONAL SPECTROSCOPY

\underline{BOGDAN NEGRU}, SCOTT J. GONCHER, AMY L. BRUNSVOLD, DANIEL M. NEUMARK, College of Chemistry, University of California, Berkeley, California 94720, and Chemical Sciences Division, Lawrence Berkeley National Laboratory, Berkeley, California 94720, USA.

Photofragment translational spectroscopy was used to study the photodissociation dynamics of the phenyl radical at 193 and 248 nm. Time of flight data collected for the C_{6}H_{4}, C_{4}H_{3}, and C_{2}H_{2} photofragments show the presence of two decomposition channels. The only C_{6}H_{5} decomposition channel observed at 248 nm corresponds to CH bond fission from the cyclic radical producing ortho-benzyne. The translational energy distribution peaks at 0 kcal/mol and is consistent with no exit barrier for the H loss process. At 193 nm photodissociation, however, H loss was observed to be the minor channel, while the major decomposition pathway corresponds with decyclization of the C_{6}H_{5} radical and subsequent fragmentation to n-C_{4}H_{3} and C_{2}H_{2}. These two momentum matched photofragments have a translational energy distribution that peaks around 9 kcal/mol, indicative of a process that proceeds through a tighter transition state. Previous theoretical work on the unimolecular decomposition of the phenyl radical^{a} predicts a second H loss process that occurs after C_{6}H_{5} decyclization resulting in the linear C_{6}H_{4} photofragment. This channel cannot be unambiguously discerned from the C_{6}H_{4}^{+} time of flight data, but is believed to take place since decyclization is observed.

aL. K. Madden, L. V. Moskaleva, S. Kristyan, and M. C. Lin J. Phys. Chem. A 1997, 101, 6790.